A car travels at 25 m/s along a horizontal curve of a radius of 450m. What is the minimum coefficient of friction between the tires and the road in order to not skid?


1
Expert's answer
2021-05-30T13:21:46-0400

As the car travels along the horizontal curve, the friction force provides necessary centripetal force:


"F_c=F_{fr},""\dfrac{mv^2}{r}=\mu mg,""\mu=\dfrac{v^2}{rg},""\mu=\dfrac{(25\ \dfrac{m}{s})^2}{450\ m\cdot9.8\ \dfrac{m}{s^2}}=0.14"